OpenStudy (anonymous):

how do i simplify each trigonometric expression. A. 1+cos^2 theta/sin^2 theta Everything Plus 1

terenzreignz (terenzreignz):

Does it look like this: \[\huge \frac{1+\cos^2(\theta)}{\sin^2(\theta)}+1\]

OpenStudy (anonymous):

change the 1 into sin^2 (theta)/sin^2 (theta), then add it to the fraction

OpenStudy (anonymous):

@terenzreignz it does!

terenzreignz (terenzreignz):

I thought so. Now follow @Peter14 's lead and simplify your way to the answer :)

OpenStudy (anonymous):

1+cos^2(θ) + sin^2(θ)/sin^2(θ)